# Spectral Truncation Kernels: Noncommutativity in $C^{*}$-algebraic Kernel Machines

### v3

@@ 8 Conclusion and Limitation
In this paper, we proposed a new class of positive definite kernels for vector- or function-valued outputs based on the spectral truncation. The proposed kernels with the framework of RKHMs resolve two shortcomings of the framework of vvRKHSs for vector- or function-valued outputs at the same time: computational cost and choice of kernels. By considering function-valued positive definite kernels, not operator-valued ones, we can alleviate the computational cost. At the same time, we can introduce noncommutativity into the learning process and can induce interactions along the data function domain. We also showed that we can control the tradeoff between the representation power and the model complexity associated with the proposed kernels. In addition, we proposed a deep approach to obtain models with higher representation powers.
@@
In the current setting, we focus on the $C^{*}$-algebra of continuous functions on the torus. Although this setting includes many important examples, e.g., periodic time-series data, continuous functions defined on a compact set in $\mathbb{R}$, finite-dimensional vectors, investigations for more general $C^{*}$-algebras allow us to apply the proposed kernel more general settings. As we discuss in Appendix D, we can generalize this setting and define positive definite kernels for more general inputs and outputs. However, theoretical investigations for these generalized settings and their computationally efficient implementations remain to be investigated in future work.
@@ Appendix A Proofs
Theorem 8    For $x,y\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$, $k_{n}^{\operatorname{prod},q}(x,y)(z)\to k^{\operatorname{prod},q}(x,y)(z)$ as $n\to\infty$. If $x$ and $y$ are differentiable, then $k_{n}^{\operatorname{sep},q}(x,y)(z)\to k^{\operatorname{sep},q}(x,y)(z)$.
Proof  We show $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$. The proofs for $k_{n}^{\operatorname{prod},q}(x,y)$ and $k_{n}^{\operatorname{sep},q}(x,y)$ are similar. We have
@@
where the sum $\sum_{j,l,r_{1},\ldots,r_{2q-1}=0}^{-1}$ is set as 00, $g_{i}(t)=\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$ for $t=[t_{1},\ldots,t_{2q}]\in\mathbb{T}^{2q}$, and $P=\{r=[r_{1},\ldots,r_{2q}]\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{m}r_{i}|\leq 1,\ l\leq m\}$. In addition, we set $j=r_{0}$ and $j=r_{q}$. The second to the last equality is derived by Lemma 9 below. Since $P$ is a convex polyhedron, $g_{i}\ast F_{n}^{2q,P}(z)\to g_{i}(z)$ as $n\to\infty$ by Lemma 10. For $k_{n}^{\operatorname{sep},q}$, we additionally use the fact $\lim_{n\to\infty}\tilde{k}(S_{n}(R_{n}(x)),S_{n}(R_{n}(y)))=\tilde{k}(\lim_{n\to\infty}S_{n}(R_{n}(x)),\lim_{n\to\infty}S_{n}(R_{n}(y)))$, which follows from the continuity of $\tilde{k}$, and the fact $S_{n}(R_{n}(x))\to x$ uniformly as $n\to\infty$ if $x$ and $y$ are differentiable (van Suijlekom, 2021, Lemma 10). $\Box$
@@
which completes the proof of the lemma. $\Box$
Proposition 12    The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep},q}$ are positive definite.
@@
for $z\in\mathbb{T}$. In addition, let $\tilde{x}_{n,j}=S_{n}(R_{n}(x_{j}))$. Then, we have
$\bigg{(}\sum_{j,l=1}^{N}d_{j}^{*}k_{n}^{\operatorname{sep},q}(x_{j},x_{l})d_{l}\bigg{)}(z)$
$=\sum_{j,l=1}^{N}\overline{d_{j}(z)}\tilde{k}(\tilde{x}_{n,j},\tilde{x}_{n,l})d_{l}(z)\sum_{j^{\prime},l^{\prime}=1}^{n}\sum_{m_{1},\ldots,m_{2q-1}=1}^{n}$
$\big{(}(R_{n}(a)^{*})\big{)}_{j^{\prime},m_{1}}\big{(}(R_{n}(a)^{*})\big{)}_{m_{1},m_{2}}\cdots\big{(}(R_{n}(a)^{*})\big{)}_{m_{q-1},m_{q}}\big{(}R_{n}(a)\big{)}_{m_{q},m_{q+1}}\cdots\big{(}R_{n}(a)\big{)}_{m_{2q-1},l^{\prime}}\mathrm{e}^{\mathrm{i}(j^{\prime}-l^{\prime})z}$
$=\sum_{j,l=1}^{N}\overline{d_{j}(z)}\tilde{k}(\tilde{x}_{n,j},\tilde{x}_{n,l})d_{l}(z)$
$\qquad\sum_{m_{q}=1}^{n}\bigg{|}\sum_{j^{\prime}=1}^{n}\sum_{m_{1},\ldots,m_{q-1}=1}^{n}\big{(}R_{n}(a)\big{)}_{m_{q},m_{1}}\cdots\big{(}R_{n}(a)\big{)}_{m_{q-2},m_{q-1}}\big{(}R_{n}(a)\big{)}_{m_{q-1},j^{\prime}}\mathrm{e}^{-\mathrm{i}j^{\prime}z}\bigg{|}^{2}$
$\geq 0.$
Thus, $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep},q}$ are positive definite. $\Box$
@@
for $z\in\mathbb{T}$. The last inequality is derived since the map $([x_{1},\ldots,x_{2q}],[y_{1},\ldots,y_{2q}])\mapsto\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x_{2j-1},y_{2j-1})}\tilde{k}_{2,j}(x_{2j},y_{2j})$ is positive definite and $F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n}\geq 0$. $\Box$
@@
Theorem 16    Assume $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep}q}$ are real-valued. Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|_{\operatorname{op}}^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$. Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$. For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
@@
Proof  We first show $(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)=\sup_{a\in\mathcal{S}}a(z)$ for $\mathcal{S}\subseteq\mathcal{A}_{+}$ for any $z\in\mathbb{T}$. Let $a\in\mathcal{S}$. Since $a(z)\leq\sup_{a\in\mathcal{S}}a(z)$ for any $z\in\mathbb{T}$, we have $a\leq_{\mathcal{A}}b$, where $b\in\mathcal{A}$ is defined as $b(z)=\sup_{a\in\mathcal{S}}a(z)$. Thus, $b$ is an upper bound of $\mathcal{S}$, and we have $\sup^{\mathcal{A}}_{a\in\mathcal{S}}a\leq_{\mathcal{A}}b$, which means $(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)\leq\sup_{a\in\mathcal{S}}a(z)$ for any $z\in\mathbb{T}$. Conversely, since $a\leq_{\mathcal{A}}\sup^{\mathcal{A}}_{a\in\mathcal{S}}a$ for $a\in\mathcal{S}$, we have $a(z)\leq(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)$ for any $z\in\mathbb{T}$. Thus, we have $\sup_{a\in\mathcal{S}}a(z)\leq(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)$.
Therefore, we have
@@
where the third equality is given by the identity $|a|_{\mathcal{A}}(z)=|a(z)|$, and the forth equality is satisfied since $\sigma_{i}$ is real-valued. $\Box$
@@
Proof  With probability at least $1-\delta$, we have
@@
$\quad\leq\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}$
$\quad=\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})(z)+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}.$
The inequality follows by Theorem 3.1 in Mohri et al. (2012). $\Box$
@@
Let $k_{n}$ be $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, or $k_{n}^{\operatorname{sep},q}$ defined in Section 3. Then, for $x\in\mathcal{A}_{0}^{d}$, $D(k_{n},x)\leq D(k_{n+1},x)$.
@@
The first inequality is derived by the following inequalities for $A\in\mathbb{C}^{n\times n}$:
@@
where $Q_{n}=[e_{1},\ldots,e_{n}]$ and $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ is the Fourier function. $\Box$
Appendix B Details of Remark 11
To understand the relationship between $q$ and the interactions, we observe the convergence shown in Lemma 10 from the perspective of $q$. The following lemma is by Brandlini and Travaglini Brandolini & Travaglini (1997).
@@ Appendix B Details of Remark 11
where $\hat{\cdot}$ is the Fourier transform, and the forth equality is by the Poisson summation formula. $\Box$
To understand the convergence, we split $F_{n}^{q,P}(t)=\sum_{m\in\mathbb{Z}^{q}}H_{n}^{q,P}(t+2\pi m)$ into two parts: $H_{n}^{q,P}(t)$ and $\tilde{H}_{n}^{q,P}(t):=\sum_{m\neq 0}H_{n}^{q,P}(t+2\pi m)$. In the convergence discussed in Lemma 10, the convolution $g\ast{H}_{n}^{q,P}(z)$ goes to $g(z)$. On the other hand, $g\ast\tilde{H}_{n}^{q,P}(z)$ goes to 00 (Brandolini & Travaglini, 1997, Theorem 1). Regarding the convergence of $g\ast{H}_{n}^{q,P}(z)$ we can evaluate it as follows, especially for $q=2$.
@@
Let ${h}^{P}_{s}(z)=\int_{t\in\mathbb{T}}\int_{sP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)$. Then, $h^{P}_{0}(z)=0$, and since $\lim_{s\to\infty}h^{P}_{s}(z)=0$, there exists $D(z)>0$ such that $|h^{P}_{s}(z)|\leq D(z)$ for any $s>0$. Thus, $|h^{P}_{jn}(z)|\leq D(z)$ for any $n\in\mathbb{N}$ and $0<j\leq 1$. Thus, by the bounded convergence theorem, we have
@@
In the same manner as $h_{s}^{P}(z)$, there exists $\tilde{C}(z)>0$ such that $|\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{1}-t)}x_{1}(t)\mathrm{d}r\mathrm{d}t|\leq\tilde{C}(z)$. By setting $C(z)=\max\{\tilde{C}(z),|x_{2}(z_{2})|\}$, we obtain the result. $\Box$
The left hand side is the average of $h_{jn}^{P}(z)$ and $h_{jn}^{Q}(z)$, where $P$ and $Q$ are symmetric with respect to the second coordinate $r_{2}$. The first term in the right hand side of Eq. (7) is described by the convergence of $h_{jn}^{[-1,1]}(z)$, which corresponds to the one-dimensional ($(q-1)$-dimensional) Fejér kernel. Since we have an additional term $1/2|\int_{\mathbb{T}^{2}}\int_{jnP_{2}}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)|$, if $C(z)>1$, the convergence of $h^{P}_{jn}(z)$ is expected to be slower than its one-dimensional counterpart $h^{[-1,1]}_{jn}(z)$. Based on this observation, we expect that the convergence in Theorem 8 becomes slower as $q$ becomes larger, and the interaction along $z\in\mathbb{T}$ becomes larger.

### v4

@@ 10 Conclusion and Discussion
In this paper, we proposed a new class of positive definite kernels for vector- or function-valued outputs based on the spectral truncation, which fills the gap between existing separable and commutative kernels. The proposed kernels with the framework of RKHMs resolve two shortcomings of the framework of vvRKHSs for vector- or function-valued outputs at the same time: computational cost and choice of kernels. By considering function-valued positive definite kernels, not operator-valued ones, we can alleviate the computational cost. At the same time, we can introduce noncommutativity into the learning process and can induce interactions along the data function domain, which enables us to control the global and local dependencies of the output on the input. In addition, we proposed a deep approach to obtain models with more flexibility.
@@
In the current setting, we focus on the $C^{*}$-algebra of continuous functions on the torus. Although this setting includes many important examples, e.g., periodic time-series data, continuous functions defined on a compact set in $\mathbb{R}$, finite-dimensional vectors, investigations for more general $C^{*}$-algebras allow us to apply the proposed kernel more general settings. As we discuss in Section 9, we can generalize this setting and define positive definite kernels for more general inputs and outputs and for more flexible frameworks. In addition, as we discuss in Section 7, there are many potential applications of the proposed kernels. Theoretical and empirical investigations for these generalized settings and applications remain to be investigated in future work.
@@ Appendix A Proofs
Theorem 10    For $x,y\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$, $k_{n}^{\operatorname{prod},q}(x,y)(z)\to k^{\operatorname{prod},q}(x,y)(z)$ as $n\to\infty$.
Proof  We show $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$. The proof for $k_{n}^{\operatorname{prod},q}(x,y)$ is similar. We have
@@
where the sum $\sum_{j,l,r_{1},\ldots,r_{2q-1}=0}^{-1}$ is set as 00, $g_{i}(t)=\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$ for $t=[t_{1},\ldots,t_{2q}]\in\mathbb{T}^{2q}$, and $P=\{r=[r_{1},\ldots,r_{2q}]\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{m}r_{i}|\leq 1,\ l\leq m\}$. In addition, we set $j=r_{0}$ and $j=r_{q}$. The second to the last equality is derived by Lemma 11 below. Since $P$ is a convex polyhedron, $g_{i}\ast F_{n}^{2q,P}(z)\to g_{i}(z)$ as $n\to\infty$ by Lemma 12.
@@
which completes the proof of the lemma.
Proposition 13    The kernel $k_{n}^{\operatorname{poly},q}$ is positive definite.
@@
for $z\in\mathbb{T}$. Thus, $k_{n}^{\operatorname{poly},q}$ is positive definite.
@@
for $z\in\mathbb{T}$. The last inequality is derived since the map $([x_{1},\ldots,x_{2q}],[y_{1},\ldots,y_{2q}])\mapsto\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x_{2j-1},y_{2j-1})}\tilde{k}_{2,j}(x_{2j},y_{2j})$ is positive definite and $F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n}\geq 0$.
@@
Theorem 17    Assume $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$ are real-valued. Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$ and $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$. Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$. For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
@@
Proof  We first show $(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)=\sup_{a\in\mathcal{S}}a(z)$ for $\mathcal{S}\subseteq\mathcal{A}_{+}$ for any $z\in\mathbb{T}$. Let $a\in\mathcal{S}$. Since $a(z)\leq\sup_{a\in\mathcal{S}}a(z)$ for any $z\in\mathbb{T}$, we have $a\leq_{\mathcal{A}}b$, where $b\in\mathcal{A}$ is defined as $b(z)=\sup_{a\in\mathcal{S}}a(z)$. Thus, $b$ is an upper bound of $\mathcal{S}$, and we have $\sup^{\mathcal{A}}_{a\in\mathcal{S}}a\leq_{\mathcal{A}}b$, which means $(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)\leq\sup_{a\in\mathcal{S}}a(z)$ for any $z\in\mathbb{T}$. Conversely, since $a\leq_{\mathcal{A}}\sup^{\mathcal{A}}_{a\in\mathcal{S}}a$ for $a\in\mathcal{S}$, we have $a(z)\leq(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)$ for any $z\in\mathbb{T}$. Thus, we have $\sup_{a\in\mathcal{S}}a(z)\leq(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)$. Therefore, we have
@@
where the third equality is given by the identity $|a|_{\mathcal{A}}(z)=|a(z)|$, and the forth equality is satisfied since $\sigma_{i}$ is real-valued.
@@
Proof  Note that we have $g(f(X_{i})(z),Y_{i}(z))=g(f(X_{i}),Y_{i})(z)$. Therefore, with probability at least $1-\delta$, we have
@@
$\quad\leq\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}.$
The inequality follows by Theorem 3.1 in Mohri et al. (2012).
@@
Let $k_{n}$ be $k_{n}^{\operatorname{poly},q}$ or $\hat{k}_{n}^{\operatorname{prod},q}$ defined in Section 3. Then, for $x\in\mathcal{A}_{0}^{d}$, $D(k_{n},x)\leq D(k_{n+1},x)$.
@@
The calculation for $\hat{k}_{n}^{\operatorname{prod},q}$ is similar. The first inequality is derived by the following inequalities for $A\in\mathbb{C}^{n\times n}$:
@@
where $Q_{n}=[e_{1},\ldots,e_{n}]$ and $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ is the Fourier function.
Appendix B Relationship between the parameter $q$ and the interactions
Theorem 10 implies that the interactions along $z\in\mathbb{T}$ in the kernels become small as $n$ grows. On the other hand, we also have the parameter $q$ that describes how many Toeplitz matrices created by $R_{n}$ are involved in the kernel. Regarding the relationship between $q$ and the interactions, we have the following remark.
Remark 26
We expect that the convergence in Theorem 10 becomes slower as $q$ becomes larger, and the interaction along $z\in\mathbb{T}$ becomes larger. This is based on the observation that the value related to the Fejér kernel on $\mathbb{T}^{q}$ is described by the corresponding value related to the Fejér kernel on $\mathbb{T}^{q-1}$, especially in the case of $q=2$.
To understand the relationship between $q$ and the interactions, we observe the convergence shown in Lemma 12 from the perspective of $q$. We have the following lemma, which is also mentioned by Brandolini and Travaglini (1997).
@@ Appendix B Relationship between the parameter $q$ and the interactions
where $\hat{\cdot}$ is the Fourier transform, and the forth equality is by the Poisson summation formula.
To understand the convergence, we split $F_{n}^{q,P}(t)=\sum_{m\in\mathbb{Z}^{q}}H_{n}^{q,P}(t+2\pi m)$ into two parts: $H_{n}^{q,P}(t)$ and $\tilde{H}_{n}^{q,P}(t):=\sum_{m\neq 0}H_{n}^{q,P}(t+2\pi m)$. In the convergence discussed in Lemma 12, the convolution $g\ast{H}_{n}^{q,P}(z)$ goes to $g(z)$. On the other hand, $g\ast\tilde{H}_{n}^{q,P}(z)$ goes to 00 (Brandolini and Travaglini, 1997, Theorem 1). Regarding the convergence of $g\ast{H}_{n}^{q,P}(z)$ we can evaluate it as follows, especially for $q=2$.
@@
Then, $h^{P}_{0}(z)=0$, and since $\lim_{s\to\infty}h^{P}_{s}(z)=0$, there exists $D(z)>0$ such that $|h^{P}_{s}(z)|\leq D(z)$ for any $s>0$. Thus, $|h^{P}_{jn}(z)|\leq D(z)$ for any $n\in\mathbb{N}$ and $0<j\leq 1$. Thus, by the bounded convergence theorem, we have
@@
In the same manner as $h_{s}^{P}(z)$, there exists $\tilde{C}(z)>0$ such that $|\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{1}-t)}x_{1}(t)\mathrm{d}r\mathrm{d}t|\leq\tilde{C}(z)$. By setting $C(z)=\max\{\tilde{C}(z),|x_{2}(z_{2})|\}$, we obtain the result.
The left-hand side of Eq. (9) is the average of $h_{jn}^{P}(z)$ and $h_{jn}^{Q}(z)$, where $P$ and $Q$ are symmetric with respect to the second coordinate $r_{2}$. The first term in the right hand side of Eq. (9) is described by the convergence of $h_{jn}^{[-1,1]}(z)$, which corresponds to the one-dimensional ($(q-1)$-dimensional) Fejér kernel. Since we have an additional term $(1/2)|\int_{\mathbb{T}^{2}}\int_{jnP_{2}}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)|$, if $C(z)>1$, the convergence of $h^{P}_{jn}(z)$ is expected to be slower than its one-dimensional counterpart $h^{[-1,1]}_{jn}(z)$. Based on this observation, we expect that the convergence in Theorem 10 becomes slower as $q$ becomes larger, and the interaction along $z\in\mathbb{T}$ becomes larger.